In lab, a student weighed out 3.281 g of an unknown salt. Her coffee cup calorimeter weighed 36.763 g. After adding water it weighed 94.732 g. Before the addition of salt, the temperature of the water in the coffee cup was 19.5degrees C d e g r e e s C . After all the salt dissolved the maximum temperature was 25degrees C d e g r e e s C . Calculate the enthalpy of solution for the salt, delta H of solution d e l t a H o f s o l u t i o n . Report your answer in J.

You can practice the math using this virtual lab. Note: Use the "Solids" tab to select a practice salt.
Answer #1

Mass of water = total mass - mass of calorimeter

= 94.732 - 36.763

= 57.969

Mass of solution = mass of water + mass of salt

= 57.969 + 3.281

= 61.25 grams

Q = mass * specific heat * ∆T

= 61.25 * 4.184 * ( 25 - 19.5)

= 1409.485 J

If you have any questions please comment

If you satisfied with the solution please rate it thanks